OpenStudy (anonymous):

please show steps Salisbury High School decided to take their students on a field trip to a theme park. A total of 150 people went on the trip. Adults pay $45.00 for a ticket and students pay $28.50 for a ticket. How many students and how many adults went to the park if they paid a total of $4770?

OpenStudy (anonymous):

45x+28.50y=4770 45*30=1350 (x=30) 28.5*120=3420 (y=120) 1350+3420=4770 45(30)+28.50(120)=4770

OpenStudy (anonymous):

Let the number of students went be x and number of adults went be y. x + y = 150 (because total 150 people went on the trip) And 28.50x + 45y = 4770 (as the question says) Eliminate x from first equation : x = 150 - y Put this value of x in second equation: 28.5(150 - y) + 45y = 4770 4275 - 28.5y + 45y = 4770 16.5y = 495 y = 30 x = 150 - y = 150 - 30 = 120 So, students were 120 and Adults were 30 that went to the trip..

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Welcome dear...

OpenStudy (anonymous):

you don't have to help Your piggy bank has a total of 46 coins in it; some are dimes and some are quarters. If you have a total of $7.00, how many quarters and how many dimes do you have?

OpenStudy (anonymous):

What is a dime and what is a quarter?? Quarter I think is 3/4 of anything... Right??

OpenStudy (anonymous):

Quarter is 25% or $0.25

OpenStudy (anonymous):

Quarter=25 cents Dime=10 cents

OpenStudy (anonymous):

dimes in $0.10

OpenStudy (anonymous):

Let there are x dimes and y quarters, then x + y = 46 or x = 46 - y 0.1x + 0.25y = 7 0.1(46 - y) + 0.25y = 7 4.6 - 0.1y + 0.25y = 7 0.15y = 2.4 y = 16 x = 46 - y = 46 - 16 = 30 So, there are 30 dimes and 16 quarters in the piggy bank...